# Know Your Mistakes: Towards Preventing Overreliance on Task-Oriented Conversational AI Through Accountability Modeling

### v3

@@ Abstract
Suvodip Dey, Yi-Jyun Sun, Gokhan Tur, Dilek Hakkani-Tür University of Illinois Urbana-Champaign sdey@illinois.edu, jennysun0830@gmail.com, {gokhan,dilek}@illinois.edu
Figure 1: Overview of Accountability Modeling. The model simultaneously generates the dialogue state and estimates slot probabilities (of being included in the dialogue state), then detects potential errors using the estimated slot probabilities. The errors can either be self-corrected or confirmed by the user via friction turns.
@@ 1 Introduction
In this work, we propose an accountability model for task-oriented conversations to mitigate user overreliance. Our approach aims to enhance the prediction of DST by detecting and correcting false positive and false negative errors, thereby ensuring greater accountability. To achieve this, we integrate an accountability head into the backbone LLMs, which is a binary classifier to predict the slots in the dialogue state. This augmented LLM not only generates the dialogue state but also estimates the probability of each slot being included in the dialogue state. The slot probabilities help to identify the possible false positive/negative slots. These errors are then corrected using a dedicated algorithm (Algo 1) that removes false positive slots and adds false negative slots (with the appropriate values), thereby improving the accuracy of DST prediction. For instance, in Fig. 1, while attraction-area is missing from the initial prediction, the accountability head assigns it a high probability of 0.81, detecting it as a potential false negative, which is subsequently corrected. Rather than self-correcting, the conversational agent can also confirm the detected errors with the user through a conversation turn, as shown in Fig. 1. In recent literature, such mindful interactions or friction turns prompting analytical thinking have been explored to address overreliance on AI Mejtoft et al. (2019); Naiseh et al. (2021); İnan et al. (2025). Therefore, slot probabilities from the accountability head can help introduce friction turns, such as confirming model uncertainty and errors, to mitigate user overreliance. The contributions of this work are as follows:
@@ 10             $B^{\prime}$[slot] $\leftarrow$ $\operatorname{generateSlotValue}(D,B,\operatorname{slot})$
The accountability head helps estimate the slot probabilities for all the slots, which can be used to detect the false positive and negative slots in the predicted dialogue state. Furthermore, the inclusion of the accountability head acts as an auxiliary loss that helps in the learning of dialogue state generation. Any information about the correct slots can make the dialogue state generation easier. Note that $\phi_{t}$ is learned to optimize both $\mathcal{L}_{\mathrm{BCE}}$ and $\mathcal{L}_{\mathrm{LM}}$. Therefore, $\phi_{t}$ encodes the knowledge about the relevant slots, which can improve the accuracy of the generation of the dialogue state.
3.3 Dialogue Sate Correction using Accountability Model
@@ 4.1 Dataset
We experiment with two widely used datasets: i) MultiWOZ Budzianowski et al. (2018), and ii) Snips Coucke et al. (2018). MultiWOZ is one of the largest multi-domain conversation corpora for task-oriented dialogue containing multi-turn conversations. We use MultiWOZ 2.4 Ye et al. (2022b), which contains fewer annotation errors and inconsistencies than the other MultiWOZ versions. On the other hand, Snips is a similar dataset for spoken language understanding, containing only single-turn conversations. The basic statistics of both datasets are shown in Table 1.
@@ 4.2 Evaluation Metric
DST is primarily evaluated using joint goal accuracy (JGA). JGA is defined as the percentage of turns where the predicted dialogue state exactly matches the ground-truth Henderson et al. (2014). Additionally, we report the Slot-F1 score, which measures the slot-level F1 performance. We also analyze false positive rate (FPR) and false negative rate (FNR) for certain results. FPR and FNR is defined as the percentage of turns containing false positive and false negative slots, respectively.
@@ 4.3 Model Architectures and Variants
We study the utility of the proposed accountability model applying it to three LLMs - Llama 3.1 (8B) Dubey and et al. (2024), Mistral (7B) Jiang et al. (2023), and Gemma (7B) Team and et al. (2024). We use the instruction-tuned version of the three models. The model nomenclature used in our experiments is described as follows.
@@ 4.4 Training Details
All the models are implemented using the PyTorch and Huggingface libraries in Python 3.12. We used LoRA Hu et al. (2021) finetuning with rank ($r$) $8$, $\alpha=32$, and dropout $0.1$. We used AdamW optimizer with learning rate 5e-5 to fine-tune both $\mathcal{M_{\operatorname{SFT}}}$ and $\mathcal{M_{\operatorname{AMD}}}$. We trained all the models for 4 epochs and chose the final model with minimum validation loss. The prompts used for model finetuning are provided in Appendix A.2. In Eqn. 4, the optimal $\lambda$ is selected based on the JGA score of the validation set. The best $\lambda$ was found to be 0.25 except for Mistral-Snips and Gemma-Snips, where $\lambda=0.1$ resulted in the best validation performance. For $\mathcal{M_{\operatorname{AMD+SC}}}$, the false positive ($\tau_{\operatorname{fp}}$) and the false negative ($\tau_{\operatorname{fn}}$) threshold is selected similarly based on the validation performance. We observed that (0.1, 0.5) and (0.05, 0.9) are the best ($\tau_{\operatorname{fp}}$, $\tau_{\operatorname{fn}}$) for MultiWOZ and Snips, respectively. These optimal values of $\lambda$, $\tau_{\operatorname{fp}}$, and $\tau_{\operatorname{fn}}$ are used to show the test performance in the rest of the article.
@@ 5.2 Impact of Varying False Positive and False Negative Threshold
Fig. 3 shows the summary of the performance gain achieved with dialogue state correction. We observe that $\mathcal{M_{\operatorname{AMD+SC}}}$ significantly enhances the performance of $\mathcal{M_{\operatorname{SFT}}}$ (9.6% relative) through self-corrections. The best performance in MultiWOZ is achieved with $\tau_{\operatorname{fp}}=0.1$ and $\tau_{\operatorname{fn}}=0.5$. However, beyond a certain point, the JGA begins to decline as false negative corrections induce false positives, as previously discussed. Figure 3 also depicts the maximum JGA achievable through dialogue state correction using $\mathcal{M_{\operatorname{AMD+OC}}}$ where OC stands for Oracle-Correction. Instead of self-correcting, $\mathcal{M_{\operatorname{AMD+OC}}}$ corrects the output of $\mathcal{M_{\operatorname{AMD}}}$ using the oracle or ground-truth slot values in the Step 2 of Algo 1. $\mathcal{M_{\operatorname{AMD+OC}}}$ does not introduce false positives while addressing false negatives, as only the slots that correspond to the ground truth are updated. The plot suggests that there is still room to achieve higher JGA through self-correction.
$\lambda$ MultiWOZ Snips
JGA Slot-F1 JGA Slot-F1
0 64.34 95.23 92.43 97.76
0.1 65.93 95.64 92.71 97.82
0.25 67.13 95.90 93.57 98.01
0.5 66.27 95.80 93.14 97.87
0.75 65.40 95.61 93.00 97.84
1.0 65.25 95.57 92.86 97.72
5.3 Impact $\lambda$ on DST performance
In this ablation study, we vary the value of $\lambda$ in Eqn. 4. In Table 6, for both datasets, we observe that the JGA and Slot-F1 increase with $\lambda$ initially and then start dropping. Note that the primary task here is to generate the dialogue state for which the JGA or Slot-F1 is estimated. Now, increasing the $\lambda$ increases the importance of the accountability head, which starts affecting the generation head after a certain point. Therefore, although any $\lambda\in[0,1]$ can improve the base model, a thorough hyperparameter optimization of $\lambda$ is required to achieve the best possible DST performance.
@@ 6 Preventing User Overeliance with Accountability Model
To evaluate this method, we conduct an experiment using a user simulator. We employ “GPT-4o mini” OpenAI and et al. (2024) as our user simulator. Given the dialogue history and a slot-value pair (detected as false negative or false positive), the user simulator is asked to confirm if it mentioned the slot-value pair by responding “Agree” (confirms interest) or “Disagree” (rejects). We correct the dialogue state by removing the false positive slots and adding the false negative slots confirmed by the user. A detailed description of the experimental setup is provided in Appendix A.2. Table 7 shows the results of this experiment where $\mathcal{M_{\operatorname{AMD+UC}}}$ denotes the User Confirmation (UC) provided by the simulator. We can observe that $\mathcal{M_{\operatorname{AMD+UC}}}$ achieves comparable performance to $\mathcal{M_{\operatorname{AMD+SC}}}$. However, further improvements could be achieved with a more advanced user simulator and refined prompt engineering. Nevertheless, the experiment highlights the utility of our proposed accountability modeling in mitigating user overreliance in real-world task-oriented conversations.
@@ 8 Limitations
  • The accountability model assumes the data to be annotated along with a fixed number of domains and slots. This is why the model cannot be directly used for new or unseen domains/slots. However, the domains/slots are generally well-defined in task-oriented conversations.
@@ Appendix A Appendix
A.1 Selection of $\lambda$
Table 8 shows the joint goal accuracy for different $\lambda$ (in Eqn. 4) on both MultiWOZ and Snips validation set. Table 8 shows the results with the Llama 3.1 model. We can observe that $\lambda=0.25$ results in the best validation performance in our setup. It also results in the best validation performance for the Mistral and Gemma on the MultiWOZ dataset. However, Mistral and Gemma show the best validation performance for the Snips dataset with $\lambda=0.1$.
@@ A.1 Selection of $\lambda$
A.2 Additional Details
This section provides additional details related to model training and prompts. The experiments are performed on Nvidia A100 machines. The hidden size of Llama and Mistral is 4096, and 3072 for Gemma. The number of trainable parameters in our LoRA fine-tuned models is 3.6M for Llama and Mistral. On the other hand, the trainable parameters of the Gemma model is 3.4M. Due to memory limitations in GPU, we use a batch size of 1. We increase the effective batch size to 8 by using gradient accumulation. The training time for MultiWOZ and Snips is approximately 20 and 6 GPU hours, respectively. All the belief states are generated using a single run.
@@ A.2 Additional Details
In Section 6, we use two kinds of prompts to confirm the predicted errors from the user-simulator. The prompt for confirming false positives is shown in Table 10, while the prompt for confirming false negatives is shown in Table 11.

### v4

@@ Abstract
Suvodip Dey, Yi-Jyun Sun, Gokhan Tur, Dilek Hakkani-Tür University of Illinois Urbana-Champaign suvodip15@gmail.com, jennysun0830@gmail.com, {gokhan,dilek}@illinois.edu
Figure 1: Overview of Accountability Modeling. The model simultaneously generates the dialogue state and estimates the probability of each slot (being included in the dialogue state). Based on these probabilities, it identifies potential errors, which can either be self-corrected using a dedicated algorithm or confirmed by the user through clarification questions.
@@ 1 Introduction
In this work, we propose an accountability model for task-oriented conversations to mitigate user overreliance. Our approach aims to enhance the prediction of DST by detecting and correcting false positive and false negative errors, thereby ensuring greater accountability. To achieve this, we integrate an accountability head into the backbone LLMs, which is a binary classifier to predict the slots in the dialogue state. This augmented LLM not only generates the dialogue state but also estimates the probability of each slot being included in the dialogue state. The slot probabilities help to identify the possible false positive/negative slots. These errors are then self-corrected using a dedicated algorithm (Algorithm 1) that removes false positive slots and adds false negative slots (with the appropriate values), thereby improving the accuracy of DST prediction. For instance in Fig. 1, the ground-truth dialogue state is {attraction-type: park, attraction-area: centre}. However, the model initially predicted {attraction-type: park}. This results in a false negative, as the prediction does not contain attraction-area: centre. The accountability head assigns a high probability (0.81) to the slot attraction-area, identifying it as a potential false negative. In Fig. 1, the model self-corrects this error by generating the value for attraction-area (i.e. centre) using Algorithm 1 and adds it to the initial prediction, thereby successfully correcting the dialogue state.
Rather than self-correcting, the conversational agent can also confirm the detected errors with the user through a conversation turn. For example in Fig. 1, the model asks a clarification question to the user to confirm the error, which helps to rectify the mistake. In recent literature, such mindful interactions or friction turns prompting analytical thinking have been explored to address overreliance on AI Mejtoft et al. (2019); Naiseh et al. (2021); İnan et al. (2025). Therefore, slot probabilities from the accountability head can help introduce friction turns, such as confirming model uncertainty and errors, to mitigate user overreliance. The contributions of this work are as follows:
@@ 10             $B^{\prime}$[slot] $\leftarrow$ $\operatorname{generateSlotValue}(D,B,\operatorname{slot})$
The accountability head helps estimate the slot probabilities for all the slots, which can be used to detect the false positive and negative slots in the predicted dialogue state. Furthermore, the inclusion of the accountability head acts as an auxiliary loss that helps in the learning of dialogue state generation. Any information about the correct slots can make the dialogue state generation easier. Note that $\phi_{t}$ is learned to optimize both $\mathcal{L}_{\mathrm{BCE}}$ and $\mathcal{L}_{\mathrm{LM}}$. Therefore, $\phi_{t}$ encodes the knowledge about the relevant slots, which can improve the accuracy of the generation of the dialogue state. A similar strategy has been shown to be effective in open-domain dialogue generation, where predicting response keywords beforehand can guide and improve response generation Dey and Desarkar (2024).
3.3 Dialogue State Correction using Accountability Model
@@ 4.1 Dataset
We experiment with two widely used datasets: i) MultiWOZ Budzianowski et al. (2018), and ii) Snips Coucke et al. (2018). MultiWOZ is one of the largest multi-domain conversation corpora for task-oriented dialogue, containing multi-turn conversations. We use MultiWOZ 2.4 Ye et al. (2022b), which contains fewer annotation errors and inconsistencies than the other MultiWOZ versions. On the other hand, Snips is a similar dataset for spoken language understanding, containing only single-turn conversations. The basic statistics of both datasets are shown in Table 1.
@@ 4.2 Evaluation Metric
DST is primarily evaluated using joint goal accuracy (JGA). JGA is defined as the percentage of turns where the predicted dialogue state exactly matches the ground-truth Henderson et al. (2014); Dey et al. (2022). Additionally, we report the Slot-F1 score, which measures the slot-level F1 performance. We also analyze false positive rate (FPR) and false negative rate (FNR) for certain results. FPR and FNR are defined as the percentage of turns containing false positive and false negative slots, respectively. Let $X$ be the number of turns containing at least one false positive slot. Let $Y$ be the number of turns containing at least one false negative slot. Then, FPR and FNR are defined as $\operatorname{FPR}=\frac{100X}{T}$ and $\operatorname{FNR}=\frac{100Y}{T}$, where $T$ is the total number of turns in the dataset.
@@ 4.3 Model Architectures and Variants
We study the utility of the proposed accountability model by applying it to three LLMs - Llama 3.1 (8B) Dubey and et al. (2024), Mistral (7B) Jiang et al. (2023), and Gemma (7B) Team and et al. (2024). We use the instruction-tuned version of the three models. The model nomenclature used in our experiments is described as follows.
@@ 4.4 Training Details
All the models are implemented using the PyTorch and Huggingface libraries in Python 3.12. We used LoRA Hu et al. (2021) finetuning with rank ($r$) $8$, $\alpha=32$, and dropout $0.1$. We used AdamW optimizer with learning rate 5e-5 to fine-tune both $\mathcal{M_{\operatorname{SFT}}}$ and $\mathcal{M_{\operatorname{AMD}}}$. We trained all the models for 4 epochs and chose the final model with the minimum validation loss. The prompts used for model finetuning are provided in Appendix A.2. In Eqn. 4, the optimal $\lambda$ is selected based on the JGA score of the validation set. The best $\lambda$ was found to be 0.25 except for Mistral-Snips and Gemma-Snips, where $\lambda=0.1$ resulted in the best validation performance. For $\mathcal{M_{\operatorname{AMD+SC}}}$, the false positive ($\tau_{\operatorname{fp}}$) and the false negative ($\tau_{\operatorname{fn}}$) threshold is selected similarly based on the validation performance. We observed that (0.1, 0.5) and (0.05, 0.9) are the best ($\tau_{\operatorname{fp}}$, $\tau_{\operatorname{fn}}$) for MultiWOZ and Snips, respectively. These optimal values of $\lambda$, $\tau_{\operatorname{fp}}$, and $\tau_{\operatorname{fn}}$ are used to show the test performance in the rest of the article.
Further details on prompts, training, and inference are provided in Section A.2.
@@ 5.2 Impact of Varying False Positive and False Negative Threshold
Fig. 3 shows the summary of the performance gain achieved with dialogue state correction. We observe that $\mathcal{M_{\operatorname{AMD+SC}}}$ significantly enhances the performance of $\mathcal{M_{\operatorname{SFT}}}$ (9.6% relative) through self-corrections. The best performance in MultiWOZ is achieved with $\tau_{\operatorname{fp}}=0.1$ and $\tau_{\operatorname{fn}}=0.5$. However, beyond a certain point, the JGA begins to decline as false negative corrections induce false positives, as previously discussed. Figure 3 also depicts the maximum JGA achievable through dialogue state correction using $\mathcal{M_{\operatorname{AMD+OC}}}$ where OC stands for Oracle-Correction. Instead of self-correcting, $\mathcal{M_{\operatorname{AMD+OC}}}$ corrects the output of $\mathcal{M_{\operatorname{AMD}}}$ using the oracle or ground-truth slot values in Step 2 of Algo 1. $\mathcal{M_{\operatorname{AMD+OC}}}$ does not introduce false positives while addressing false negatives, as only the slots that correspond to the ground truth are updated. The plot suggests that there is still room to achieve higher JGA through self-correction.
$\lambda$ MultiWOZ Snips
JGA Slot-F1 JGA Slot-F1
0 64.34 95.23 92.43 97.76
0.1 65.93 95.64 92.71 97.82
0.25 67.13 95.90 93.57 98.01
0.5 66.27 95.80 93.14 97.87
0.75 65.40 95.61 93.00 97.84
1.0 65.25 95.57 92.86 97.72
5.3 Impact $\lambda$ on DST performance
In this ablation study, we vary the value of $\lambda$ in Eqn. 4. In Table 5, for both datasets, we observe that the JGA and Slot-F1 increase with $\lambda$ initially and then start dropping. Note that the primary task here is to generate the dialogue state for which the JGA or Slot-F1 is estimated. Now, increasing the $\lambda$ increases the importance of the accountability head, which starts affecting the generation head after a certain point. Therefore, although any $\lambda\in[0,1]$ can improve the base model, a thorough hyperparameter optimization of $\lambda$ is required to achieve the best possible DST performance. The validation performance used to select the optimal value of $\lambda$ is discussed in Appendix A.1.
@@ 6 Preventing User Overeliance with Accountability Model
To evaluate this method, we conduct an experiment using a user simulator. We employ “GPT-4o mini” OpenAI and et al. (2024) as our user simulator, assuming cooperative user behavior. Given the dialogue history and a slot-value pair (detected as false negative or false positive), the user simulator is asked to confirm if it mentioned the slot-value pair by responding “Agree” (confirms interest) or “Disagree” (rejects). We correct the dialogue state by removing the false positive slots and adding the false negative slots confirmed by the user. A detailed description of the experimental setup is provided in Appendix A.2.
Table 7 shows the results of this experiment where $\mathcal{M_{\operatorname{AMD+UC}}}$ denotes the User Confirmation (UC) provided by the simulator. We can observe that $\mathcal{M_{\operatorname{AMD+UC}}}$ achieves comparable performance to $\mathcal{M_{\operatorname{AMD+SC}}}$. Although the user simulator may not fully reflect actual user behavior, assuming a cooperative user allows us to interpret the results in Table 7 as a lower bound on achievable DST performance. A real cooperative user would likely be even more effective in identifying and confirming errors, potentially leading to further performance improvements. Additional improvements may also be possible through the use of more advanced simulators and refined prompt engineering. Nonetheless, this experiment demonstrates the practical value of our accountability modeling framework in reducing user overreliance in real-world task-oriented dialogues.
@@ 8 Limitations
  • The accountability model assumes the data to be annotated along with a fixed number of domains and slots. This is why the proposed model cannot be directly used for new or unseen domains/slots. However, in task-oriented conversations, domains and slots are typically well-defined, making the model effectively trainable. Furthermore, the accountability model outperforms few-shot and zero-shot methods, as shown in Table 3. While general LLMs (such as GPT-4o) can operate with limited data, their performance remains suboptimal. In task-oriented conversations, accurately understanding intent is essential for effective dialogue management. Moreover, disregarding the training overhead, a small LLM with accountability modeling is more cost-effective than a large general-purpose LLM for dialogue state tracking in real applications.
@@ Appendix A Appendix
A.1 Selection of $\lambda$
Table 8 shows the joint goal accuracy for different $\lambda$ (in Eqn. 4) on both the MultiWOZ and Snips validation sets. Table 8 shows the results with the Llama 3.1 model. We can observe that $\lambda=0.25$ results in the best validation performance in our setup. It also results in the best validation performance for Mistral and Gemma on the MultiWOZ dataset. However, Mistral and Gemma show the best validation performance for the Snips dataset with $\lambda=0.1$.
A.2 Additional Details
This section provides additional details related to model training, inference, and prompts. The experiments are performed on Nvidia A100 machines. The hidden size of Llama and Mistral is 4096 and 3072 for Gemma. The number of trainable parameters in our LoRA fine-tuned models is 3.6M for Llama and Mistral. On the other hand, the trainable parameters of the Gemma model is 3.4M. We found that 8B models provided an optimal balance for training on our GPU servers. Smaller models resulted in lower performance, while larger models posed computational challenges due to resource constraints. This is why we conducted our experiments using LLaMA 3.1 (8B), Mistral (7B), and Gemma (7B). It also helped to balance between performance and computational efficiency, making our approach both scalable and practical for real-world applications. Due to memory limitations in GPU, we use a batch size of 1. We increase the effective batch size to 8 by using gradient accumulation. The training time for MultiWOZ and Snips is approximately 20 and 6 GPU hours, respectively.
During inference, the model takes only the dialogue context as input and generates the corresponding dialogue state, without relying on any oracle belief state. We do not apply any additional post-processing to enforce JSON formatting. In our experiments, all outputs adhered to the expected JSON format, and no formatting issues were observed. However, in the event of a failure, the output will default to an empty dictionary (i.e., {}). All the belief states are generated using a single run.
In Section 6, we used two kinds of prompts to confirm the predicted errors from the user-simulator. The prompt for confirming false positives is shown in Table 10, while the prompt for confirming false negatives is shown in Table 11.